OpenStudy (anonymous):

Solve the equation using inverse operations. Check your solutions. In your final answer, include all of your work. 2 + x^2 = 18

OpenStudy (anonymous):

x=9

OpenStudy (kiakat):

You're first step is to isolate the variable. How would you do that?

OpenStudy (anonymous):

i dont know .. can u show me how to do it @KiaKat

OpenStudy (kiakat):

Isolating the variable means you have to get the x term on one side. What do you have move in order to do that??

OpenStudy (anonymous):

i dont understand sorry

OpenStudy (kiakat):

okay well you need to get x^2 on one side of the equation. So to do that you'd have to subtract the +2 from both sides of the equation. 2+x^2=18 -2 -2 so then you'd have x^2=16

OpenStudy (anonymous):

oh okay thanks